OpenStudy (anonymous):

Which graph represents a phase shift of pi/2 units right for the graph of y=sinx?

OpenStudy (anonymous):

You have no pictures of graphs

OpenStudy (inkyvoyd):

that would be -cos(x)

OpenStudy (anonymous):

y=sin(x-pi/2) y=secx cscx cosx

OpenStudy (inkyvoyd):

Well.

OpenStudy (inkyvoyd):

Since you obviously can't do trig identities, let's have some fun!

OpenStudy (inkyvoyd):

sin(-x)=-sin(x)

OpenStudy (inkyvoyd):

sin(x+y)=sinxcosy+sinycosx

OpenStudy (inkyvoyd):

Now, x=x, -pi/2=y

OpenStudy (inkyvoyd):

also, cos(-x)=cos(x)

OpenStudy (inkyvoyd):

ok. sin(x-pi/2) sin(x-90) -> convert to degrees because i'm a noob sin(x-90)=sinxcos(-90)+sin(-90)cos(x) =sinx*0-1*cos(x) =-cos(x)

OpenStudy (inkyvoyd):

Did you see what I did there? I gave youthe answer, and showed you why my answer was the same.

OpenStudy (anonymous):

thank u so much!!!

OpenStudy (inkyvoyd):

But, in general, a phase shift in a graph is represented by sin(x-theta), or cos(x-theta).

OpenStudy (anonymous):

Which of the following is a sinusoid? a) y= [x] b) y = 3 sqr x c) x^2 + y^2 = 1 d) y=cosx

OpenStudy (anonymous):

The answer to the second question was D.